# Rethinking Debiasing: Real-World Bias Analysis and Mitigation

### v3

@@ C.2 Proof of proposition 1
Proposition 1. Assume feature $y^{s}=+1$ is biased. Then high bias prevalence distribution, i.e. feature $y^{s}=-1$ is biased as well, implying that the marginal distribution of $y^{t}$ and $y^{s}$ is matched, i.e. $p^{t}_{+}=p^{s}_{+}$. Specifically, as $\theta$ approaches to $1$, the marginal distribution of $y^{s}$ approaches to that of $y^{t}$, i.e. $lim_{\theta\to 1}p^{s}_{+}=p^{t}_{+}$.
Proof. We first derive the upper and lower bound of the $p^{s}_{+}$, and then we can prove the proposition with the squeeze theorem [41].
@@
$\rho_{+}>\theta\cdot\rho_{max}^{+}=\theta\cdot(1-p^{t}_{+}),$ (9)
$\rho_{-}>\theta\cdot\rho_{max}^{-}=\theta\cdot p^{t}_{+}$ (10)
@@
We can also derive the simplified bias magnitude of feature $y^{s}=-1$ based on the conditional distribution, and find its relationship with $\rho_{+}$:
$\rho_{-}$ $=\tau_{-}-p^{t}_{-}$ (11)
$=\frac{1-p^{t}_{+}-p^{s}_{+}(1-\tau_{+})}{1-p^{s}_{+}}-(1-p^{t}_{+})$ (12)
$=\frac{p^{s}_{+}(\tau_{+}-p^{t}_{+})}{1-p^{s}_{+}}$ (13)
$=\frac{p^{s}_{+}}{1-p^{s}_{+}}\rho_{+}$ (14)
We can then derive the lower bound of $p^{s}_{+}$ with the above equation and inequalities:
$\frac{p^{s}_{+}}{1-p^{s}_{+}}(1-p^{t}_{+})\geq\frac{p^{s}_{+}}{1-p^{s}_{+}}\rho_{+}=\rho_{-}\geq\theta\cdot p^{t}_{+}$ (15)
$p^{s}_{+}\geq\frac{\theta\cdot p^{t}_{+}}{1-p^{t}_{+}+\theta\cdot p^{t}_{+}}\geq\theta\cdot p^{t}_{+}=LB(\theta)$ (16)
We can also derive the following equation and inequalities of $\tau_{+}$ according to its definition.
$\tau_{+}=\frac{p^{s}_{+}\cdot P(y^{s}=+1|y^{t}=+1)}{p^{s}_{+}}\leq\frac{p^{t}_{+}}{p^{s}_{+}}$ (17)
$\tau_{+}=p^{t}_{+}+\rho_{+}\geq\theta(1-p^{t}_{+})+p^{t}_{+}$ (18)
Then we can derive the upper bound of $p^{s}_{+}$:
$\theta(1-p^{t}_{+})+p^{t}_{+}\leq\tau_{+}\leq\frac{p^{t}_{+}}{p^{s}_{+}}$ (19)
$p^{s}_{+}\leq\frac{p^{t}_{+}}{\theta(1-p^{t}_{+})+p^{t}_{+}}=UB(\theta)$ (20)
@@
$\lim_{\theta\to 1}LB(\theta)=\lim_{\theta\to 1}\theta\cdot p^{t}_{+}=p^{t}_{+}$ (21)
$\lim_{\theta\to 1}UB(\theta)=\lim_{\theta\to 1}\frac{p^{t}_{+}}{\theta(1-p^{t}_{+})+p^{t}_{+}}=p^{t}_{+}$ (22)
@@
$LB(\theta)\leq p^{s}_{+}\leq UB(\theta)$ (23)
$\lim_{\theta\to 1}p^{s}_{+}=\lim_{\theta\to 1}LB(\theta)=\lim_{\theta\to 1}UB(\theta)=p^{t}_{+}$ (24)
@@ C.3 Proof of proposition 2
Proposition 2. Given that the marginal distribution of $y^{s}$ and $y^{t}$ are matched and not uniform, i.e. $p=p^{s}_{+}=p^{t}_{+}<0.5$. The bias magnitude of sparse feature, i.e. $\rho^{*}_{+}$, is monotone decreasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{+}=-log(1-\phi_{+})$.The bias magnitude of the dense feature, i.e. $\rho^{*}_{-}$, is monotone increasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{-}=0$.
Proof. Given the distribution proposed in section C.1 and the condition $p=p^{s}_{+}=p^{t}_{+}<0.5$, we further use $\phi_{+}=\frac{\rho_{+}}{\rho^{max}_{+}}$ to express $\tau$:
$\tau_{+}=p+\phi_{+}(1-p)$ (25)
$\tau_{-}=1-p+\phi_{+}\cdot p$ (26)
We can then derive the bias magnitude of the sparse feature $y^{s}=+1$, given $p=p^{s}_{+}=p^{t}_{+}<0.5$, and warp it with a function $t(p)$.
$\rho^{*}_{+}$ $=KL(P(y^{t}),P(y^{t}|y^{s}=+1))$ (27)
$=p\cdot log(\frac{p}{\tau_{+}})+(1-p)\cdot log(\frac{1-p}{1-\tau_{+}})$ (28)
$=p\cdot log(\frac{p}{p+\phi_{+}(1-p)})+(1-p)\cdot log(\frac{1-p}{1-p-\phi_{+}(1-p)})$ (29)
$=p\cdot log(\frac{p}{p+\phi_{+}(1-p)})+(1-p)\cdot log(\frac{1}{1-\phi_{+}})$ (30)
$=p\cdot log(\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)})+log(\frac{1}{1-\phi_{+}})=t(p)$ (31)
We further derive the partial derivative of $\rho^{*}_{+}$ on $p$ as follows:
$\frac{\partial t(p)}{\partial p}$ $=p\cdot log(\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)})+1-\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)}$ (32)
Here we apply substitution method to replace $\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)}$ with $x$:
@@
Thus, we have $f(x)<0$ in the interval $0<x\leq 1$, proving $\rho^{*}_{+}=t(p)$ to be monotone decreasing at $p$.
$\frac{\partial\rho^{*}_{+}}{\partial p}=\frac{\partial t(p)}{\partial p}<0$ (37)
Similarly, we can derive the bias magnitude of the dense feature $y^{s}=-1$, and see that it is just $t(1-p)$
$\rho^{*}_{-}$ $=KL(P(y^{t}),P(y^{t}|y^{s}=-1))$ (38)
$=(1-p)\cdot log(\frac{(1-p)(1-\phi_{+})}{1-p+\phi_{+}\cdot p})+log(\frac{1}{1-\phi_{+}})$ (39)
@@
As a result, we can prove the monotonicity of $\rho^{*}_{-}$ with the chain rule.
$\frac{\partial\rho^{*}_{-}}{\partial p}$ $=\frac{\partial t(1-p)}{\partial p}$ (41)
@@
We can then derive the convergence of sparse feature bias magnitude $\rho^{*}_{+}$ when $p$ approaches 0 with L’Hôpital’s Rule [41].
$\lim_{p\to 0^{+}}\rho^{*}_{+}$ $=\lim_{p\to 0^{+}}t(p)$ (45)
$=\lim_{p\to 0^{+}}(p\cdot log(\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)}))+log(\frac{1}{1-\phi_{+}})$ (46)
$=\lim_{p\to 0^{+}}(p\cdot log(p))+\lim_{p\to 0^{+}}(p\cdot log(\frac{1-\phi_{+}}{p+\phi_{+}(1-p)}))+log(\frac{1}{1-\phi_{+}})$ (47)
$=\lim_{p\to 0^{+}}\frac{log(p)}{\frac{1}{p}}+log(\frac{1}{1-\phi_{+}})$ (48)
$=\lim_{p\to 0^{+}}\frac{(log(p))^{\prime}}{(\frac{1}{p})^{\prime}}+log(\frac{1}{1-\phi_{+}})$ (49)
$=\lim_{p\to 0^{+}}\frac{\frac{1}{p}}{-\frac{1}{p^{2}}}+log(\frac{1}{1-\phi_{+}})$ (50)
$=log(\frac{1}{1-\phi_{+}})$ (51)
Similarly, we can derive the convergence of dense feature bias magnitude $\rho^{*}_{-}$ when $p$ approaches to 0.
$\lim_{p\to 0^{+}}\rho^{*}_{-}$ $=\lim_{p\to 0^{+}}t(1-p)$ (52)
$=\lim_{p\to 1^{-}}(p\cdot log(\frac{p(1-\phi_{+})}{p+\phi_{+}(1-p)}))+log(\frac{1}{1-\phi_{+}})$ (53)
$=log(1-\phi_{+})+log(\frac{1}{1-\phi_{+}})$ (54)
@@ D.3 Baselines
JTT [16].
JTT is a classic DBAM method that has been applied to both the image and NLP domains. JTT identifies challenging examples by training an initial model using standard empirical risk minimization (ERM) and collecting misclassified examples into an error set. The second stage involves re-training the model while upweighting the error set to prioritize examples that the first-stage model struggled with. This approach aims to address performance disparities caused by spurious correlations, leading to better generalization across groups with minimal additional annotation costs.
@@ D.5 Design of feature destructing methods
D.6 Applying DiD to DBAM methods
As aforementioned in the main paper, when applying our method to the existing DBAM methods [12, 9, 20], we do not modify the training procedure of the debiased model $M_{d}$. For both methods, we train the biased model $M_{b}$ with target feature destroyed data. This is done by simply adding a feature destructive data transformation during data processing, with minimal computational overhead.
@@ E.1 Detailed results and explanations of the main experiments
The main results in the main paper are presented in the form of performance gain and only contain results of BC accuracy and average accuracy on the unbiased test set, here we present the results in their original form, together with error bars, detailed results of accuracies for BA and BN samples of each dataset as well. Results on the Colored MNIST and Corrupted CIFAR10 datasets can be found in Table 6 and Table 7, respectively. It shows that combining DiD not only boosts the performance of existing DBAM methods but also achieves the best performances.
The performance generally varies between different datasets, different types of biased distribution, and algorithms with and without BiasEnsemble, e.g. between LfF and BE LfF. Firstly, the inconsistency between datasets is likely to depend on how thoroughly the target feature is destroyed within the dataset. The target features of Colored MNIST, i.e. digits, are destroyed more completely by patch shuffling, for shape is the only feature within digits. In comparison, the target feature of Corrupted CIFAR10 is more complicated (including shape, texture, color, etc.), and thus can not be thoroughly destroyed by patch shuffling, causing relatively lower performance gain. Secondly, the performance inconsistency between different biased distributions is due to the reliance of existing DBAM methods on the high bias prevalence assumption for bias capturing as discussed in section 4.2. Specifically, as the bias prevalence of the training distribution becomes higher, better bias capture can be achieved by existing DBAM even without our method, thus making our improvement on the performance less significant. This conclusion is supported by our experimental results shown in Figure 5. As for the performance inconsistency between algorithms with and without BiasEnsemble, it is due to the fact that BiasEnsemble is also a method targeted to enhance the bias capture procedure of the debiasing framework. As we can see that BiasEnsemble is much more robust to the change in the bias magnitude and prevalence from Table 1. In other words, certain overlap between the goals of BiasEnsemble and our method resulted in smaller improvement of our method on BiasEnsemble-based baselines.
Table 6: Results on Colored MNIST dataset show that combining DiD not only boosts the performance of existing DBAM methods but also achieves the best performances. The accuracy of BN samples is marked as ’-’ in LMLP and HMHP distribution for there is no BN sample within the dataset according to our evaluation setting in Appendix D.
@@
Table 7: Results on Corrupted CIFAR10 dataset show that combining DiD not only boosts the performance of existing DBAM methods but also achieves the best performances. The accuracy of BN samples is marked as ’-’ in LMLP and HMHP distribution for there is no BN sample within the dataset according to our evaluation setting in Appendix D.
Distr. Algorithm Accuracy BA acc BC acc BN acc Avg acc LMLP ERM 80.40 ± 0.81 62.50 ± 0.15 - 64.29 ± 0.06 LfF 59.13 ± 0.68 55.03 ± 0.04 - 55.44 ± 0.09 + DiD 69.47 ± 0.96 62.04 ± 0.21 - 62.78 ± 0.10 BE LfF 70.87 ± 1.30 52.10 ± 0.30 - 53.98 ± 0.40 + DiD 63.23 ± 2.10 53.21 ± 0.20 - 54.21 ± 0.38 DisEnt 61.58 ± 0.57 55.45 ± 0.23 - 56.06 ± 0.17 + DiD 72.23 ± 0.74 60.84 ± 0.40 - 61.98 ± 0.30 BE DisEnt 62.73 ± 0.61 56.59 ± 0.08 - 57.20 ± 0.13 + DiD 65.98 ± 0.40 60.92 ± 0.20 - 61.42 ± 0.21 HMLP ERM 84.67 ± 0.64 55.85 ± 0.17 65.75 ± 0.00 65.05 ± 0.13 LfF 73.33 ± 1.67 47.70 ± 0.58 54.58 ± 0.49 54.15 ± 0.41 + DiD 78.67 ± 2.14 54.81 ± 2.26 63.71 ± 2.69 63.06 ± 2.63 BE LfF 70.33 ± 2.19 50.96 ± 2.35 54.14 ± 0.25 54.02 ± 0.36 + DiD 68.80 ± 0.88 50.20 ± 0.79 54.39 ± 0.18 54.15 ± 0.15 DisEnt 61.67 ± 1.67 52.48 ± 0.56 54.65 ± 0.56 54.53 ± 0.49 + DiD 73.67 ± 2.64 55.26 ± 0.93 62.11 ± 0.17 61.61 ± 0.13 BE DisEnt 75.33 ± 5.21 49.15 ± 1.54 56.86 ± 0.30 56.35 ± 0.35 + DiD 78.40 ± 1.00 54.09 ± 1.07 62.05 ± 0.34 61.50 ± 0.38 HMHP ERM 89.97 ± 0.34 29.37 ± 0.30 - 35.43 ± 0.24 LfF 72.70 ± 0.81 35.30 ± 0.33 - 39.04 ± 0.33 + DiD 82.07 ± 1.09 37.05 ± 0.31 - 41.55 ± 0.19 BE LfF 82.73 ± 0.92 31.48 ± 0.82 - 36.61 ± 0.65 + DiD 78.30 ± 0.47 32.90 ± 1.79 - 37.44 ± 1.61 DisEnt 70.77 ± 2.27 36.04 ± 0.62 - 39.51 ± 0.36 + DiD 76.60 ± 0.70 39.05 ± 0.35 - 42.80 ± 0.25 BE DisEnt 78.60 ± 1.56 34.20 ± 0.43 - 38.64 ± 0.38 + DiD 78.70 ± 1.47 37.72 ± 0.96 - 41.82 ± 0.91
E.2 Hyper-parameter sensitivity
As shown in Table 9, we examine three feature destruction methods: pixel-shuffling, patch-shuffling, and center occlusion, to destroy object shapes. We observed that patch-shuffle with patch-size 8 exhibits the best performance on Corrupted CIFAR10 which is of size 32x32.
E.3 Application of DiD on the bias detection task

### v4

@@ C.2 Proof of proposition 1
Proposition 1. Assume feature $y^{s}=a_{j}^{s}$ is biased. Then high bias prevalence distribution, i.e. feature $y^{s}\neq a_{j}^{s}$ is biased as well, implies that the marginal distribution of $a_{i}^{t}$ and $a_{j}^{s}$ is matched, i.e. $lim_{\theta\to 1}p^{s}_{j}=p^{t}_{i}$.
Proof. We first derive the upper and lower bound of the $p^{s}_{j}$, and then we can prove the proposition with the squeeze theorem [45].
@@
$\rho_{j}>\theta\cdot\rho^{max}_{j}=\theta\cdot(1-p^{t}_{i}),$ (10)
$\rho_{-}>\theta\cdot\rho^{max}_{\neq j}=\theta\cdot p^{t}_{i}$ (11)
@@
We can also derive the simplified bias magnitude of feature $y^{s}\neq a_{j}^{s}$ based on the conditional distribution, and find its relationship with $\rho_{j}$:
$\rho_{-}$ $=\tau_{\neq j}-p^{t}_{-}$ (12)
$=\frac{1-p^{t}_{i}-p^{s}_{j}(1-\tau_{j})}{1-p^{s}_{j}}-(1-p^{t}_{i})$ (13)
$=\frac{p^{s}_{j}(\tau_{j}-p^{t}_{i})}{1-p^{s}_{j}}$ (14)
$=\frac{p^{s}_{j}}{1-p^{s}_{j}}\rho_{j}$ (15)
We can then derive the lower bound of $p^{s}_{j}$ with the above equation and inequalities:
$\frac{p^{s}_{j}}{1-p^{s}_{j}}(1-p^{t}_{i})\geq\frac{p^{s}_{j}}{1-p^{s}_{j}}\rho_{j}=\rho_{-}\geq\theta\cdot p^{t}_{i}$ (16)
$p^{s}_{j}\geq\frac{\theta\cdot p^{t}_{i}}{1-p^{t}_{i}+\theta\cdot p^{t}_{i}}\geq\theta\cdot p^{t}_{i}=LB(\theta)$ (17)
We can also derive the following equation and inequalities of $\tau_{j}$ according to its definition.
$\tau_{j}=\frac{p^{s}_{j}\cdot P(y^{s}=a_{j}^{s}|y^{t}=a_{i}^{t})}{p^{s}_{j}}\leq\frac{p^{t}_{i}}{p^{s}_{j}}$ (18)
$\tau_{j}=p^{t}_{i}+\rho_{j}\geq\theta(1-p^{t}_{i})+p^{t}_{i}$ (19)
Then we can derive the upper bound of $p^{s}_{j}$:
$\theta(1-p^{t}_{i})+p^{t}_{i}\leq\tau_{j}\leq\frac{p^{t}_{i}}{p^{s}_{j}}$ (20)
$p^{s}_{j}\leq\frac{p^{t}_{i}}{\theta(1-p^{t}_{i})+p^{t}_{i}}=UB(\theta)$ (21)
@@
$\lim_{\theta\to 1}LB(\theta)=\lim_{\theta\to 1}\theta\cdot p^{t}_{i}=p^{t}_{i}$ (22)
$\lim_{\theta\to 1}UB(\theta)=\lim_{\theta\to 1}\frac{p^{t}_{i}}{\theta(1-p^{t}_{i})+p^{t}_{i}}=p^{t}_{i}$ (23)
@@
$LB(\theta)\leq p^{s}_{j}\leq UB(\theta)$ (24)
$\lim_{\theta\to 1}p^{s}_{j}=\lim_{\theta\to 1}LB(\theta)=\lim_{\theta\to 1}UB(\theta)=p^{t}_{i}$ (25)
@@ C.3 Proof of proposition 2
Proposition 2. Given that the marginal distribution of $a_{j}^{s}$ and $a_{i}^{t}$ are matched and not uniform, i.e. $p=p^{s}_{i}=p^{t}_{j}<0.5$. The bias magnitude of sparse feature, i.e. $\rho^{*}_{j}$, is monotone decreasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{j}=-log(1-\phi_{j})$. The bias magnitude of the other features, i.e. $\rho^{*}_{\neq j}$, is monotone increasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{\neq j}=0$.
Proof. Given the distribution proposed in section C.1 and the condition $p=p^{s}_{j}=p^{t}_{i}<0.5$, we further use $\phi_{j}=\frac{\rho_{j}}{\rho^{max}_{j}}$ to express $\tau$:
$\tau_{j}=p+\phi_{j}(1-p)$ (26)
$\tau_{\neq j}=1-p+\phi_{j}\cdot p$ (27)
We can then derive the bias magnitude of the sparse feature $y^{s}=a_{j}^{s}$, given $p=p^{s}_{j}=p^{t}_{i}<0.5$, and warp it with a function $t(p)$.
$\rho^{*}_{j}$ $=KL(P(y^{t}),P(y^{t}|y^{s}=a_{j}^{s}))$ (28)
$=p\cdot log(\frac{p}{\tau_{j}})+(1-p)\cdot log(\frac{1-p}{1-\tau_{j}})$ (29)
$=p\cdot log(\frac{p}{p+\phi_{j}(1-p)})+(1-p)\cdot log(\frac{1-p}{1-p-\phi_{j}(1-p)})$ (30)
$=p\cdot log(\frac{p}{p+\phi_{j}(1-p)})+(1-p)\cdot log(\frac{1}{1-\phi_{j}})$ (31)
$=p\cdot log(\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)})+log(\frac{1}{1-\phi_{j}})=t(p)$ (32)
We further derive the partial derivative of $\rho^{*}_{j}$ on $p$ as follows:
$\frac{\partial t(p)}{\partial p}$ $=p\cdot log(\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)})+1-\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)}$ (33)
Here we apply substitution method to replace $\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)}$ with $x$:
@@
Thus, we have $f(x)<0$ in the interval $0<x\leq 1$, proving $\rho^{*}_{j}=t(p)$ to be monotone decreasing at $p$.
$\frac{\partial\rho^{*}_{j}}{\partial p}=\frac{\partial t(p)}{\partial p}<0$ (38)
Similarly, we can derive the bias magnitude of the dense feature $y^{s}\neq a_{j}^{s}$, and see that it is just $t(1-p)$
$\rho^{*}_{\neq j}$ $=KL(P(y^{t}),P(y^{t}|y^{s}\neq a_{j}^{s}))$ (39)
$=(1-p)\cdot log(\frac{(1-p)(1-\phi_{j})}{1-p+\phi_{j}\cdot p})+log(\frac{1}{1-\phi_{j}})$ (40)
@@
As a result, we can prove the monotonicity of $\rho^{*}_{\neq j}$ with the chain rule.
$\frac{\partial\rho^{*}_{\neq j}}{\partial p}$ $=\frac{\partial t(1-p)}{\partial p}$ (42)
@@
We can then derive the convergence of sparse feature bias magnitude $\rho^{*}_{j}$ when $p$ approaches 0 with L’Hôpital’s Rule [45].
$\lim_{p\to 0^{+}}\rho^{*}_{j}$ $=\lim_{p\to 0^{+}}t(p)$ (46)
$=\lim_{p\to 0^{+}}(p\cdot log(\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)}))+log(\frac{1}{1-\phi_{j}})$ (47)
$=\lim_{p\to 0^{+}}(p\cdot log(p))+\lim_{p\to 0^{+}}(p\cdot log(\frac{1-\phi_{j}}{p+\phi_{j}(1-p)}))+log(\frac{1}{1-\phi_{j}})$ (48)
$=\lim_{p\to 0^{+}}\frac{log(p)}{\frac{1}{p}}+log(\frac{1}{1-\phi_{j}})$ (49)
$=\lim_{p\to 0^{+}}\frac{(log(p))^{\prime}}{(\frac{1}{p})^{\prime}}+log(\frac{1}{1-\phi_{j}})$ (50)
$=\lim_{p\to 0^{+}}\frac{\frac{1}{p}}{-\frac{1}{p^{2}}}+log(\frac{1}{1-\phi_{j}})$ (51)
$=log(\frac{1}{1-\phi_{j}})$ (52)
Similarly, we can derive the convergence of dense feature bias magnitude $\rho^{*}_{\neq j}$ when $p$ approaches to 0.
$\lim_{p\to 0^{+}}\rho^{*}_{\neq j}$ $=\lim_{p\to 0^{+}}t(1-p)$ (53)
$=\lim_{p\to 1^{-}}(p\cdot log(\frac{p(1-\phi_{j})}{p+\phi_{j}(1-p)}))+log(\frac{1}{1-\phi_{j}})$ (54)
$=log(1-\phi_{j})+log(\frac{1}{1-\phi_{j}})$ (55)
@@ D.3 Baselines
JTT [16].
JTT is a classic debiasing method w/o bias label, that has been applied to both the image and NLP domains. JTT identifies challenging examples by training an initial model using standard empirical risk minimization (ERM) and collecting misclassified examples into an error set. The second stage involves re-training the model while upweighting the error set to prioritize examples that the first-stage model struggled with. This approach aims to address performance disparities caused by spurious correlations, leading to better generalization across groups with minimal additional annotation costs.
@@
DPR [20].
DPR is another recently proposed debiasing method w/o bias label. DPR rectifies biased models through fine-tuning. They construct a small pivotal subset with a higher proportion of bias-conflicting samples using BCSI, which serves as an effective alternative to an unbiased set. Leveraging this pivotal set, they rectify a biased model through fine-tuning with only a few additional iterations.
DeNetDM [19]
. DeNetDM is another recently proposed debiasing method w/o bias label. utilize a technique inspired by the Product of Experts, where one expert is deeper than the other. They propose a strategy where they train a deep debiased model utilizing the information acquired from both deep (perfectly biased) and shallow (weak debiased) network in the previous phase.
@@ D.5 Design of feature destructing methods
D.6 Applying DiD to existing methods
As aforementioned in the main paper, when applying our method to the existing debiasing methods [12, 9, 23], we do not modify the training procedure of the debiased model $M_{d}$. For both methods, we train the biased model $M_{b}$ with target feature destroyed data. This is done by simply adding a feature destructive data transformation during data processing, with minimal computational overhead.
@@ Appendix E Additional Empirical Results
(a) DisEnt
(b) LfF
(c) Colored MNIST
(d) Corrupted CIFAR10
Figure 7: Figure 7(a) and 7(b) support our claim in section 4 that existing debiasing methods tend to overlook BN samples when training on LP distributions. Figure 7(a) and 7(b) show that our approach effectively emphasizes BN samples by raising its weights.
@@ E.1 Detailed results and explanations of the main experiments
The main results in the main paper are presented in the form of performance gain and only contain results of BC accuracy and average accuracy on the unbiased test set, here we present the results in their original form, together with error bars, detailed results of accuracies for BA and BN samples of each dataset as well. Results on the Colored MNIST and Corrupted CIFAR10 datasets can be found in Table 5 and Table 6, respectively. It shows that combining DiD not only boosts the performance of existing debiasing methods but also achieves the best performance.
The performance generally varies between different datasets, different types of biased distribution, and algorithms with and without BiasEnsemble, e.g. between LfF and BEL. Firstly, the inconsistency between datasets is likely to depend on how thoroughly the target feature is destroyed within the dataset. The target features of Colored MNIST, i.e. digits, are destroyed more completely by patch shuffling, for shape is the only feature within digits. In comparison, the target feature of Corrupted CIFAR10 is more complicated (including shape, texture, color, etc.), and thus can not be thoroughly destroyed by patch shuffling, causing relatively lower performance gain. Secondly, the performance inconsistency between different biased distributions is due to the reliance of existing debiasing methods on the high bias prevalence assumption for bias capturing as discussed in section 4.2. Specifically, as the bias prevalence of the training distribution becomes higher, better bias capture can be achieved even without our method, thus making our improvement on the performance less significant, but still quite effective. This conclusion is supported by our experimental results shown in Figure 5. As for the performance inconsistency between algorithms with and without BiasEnsemble, it is due to the fact that BiasEnsemble is also a method targeted to enhance the bias capture procedure of the debiasing framework. As we can see that BiasEnsemble is much more robust to the change in the bias magnitude and prevalence from Table 1. In other words, certain overlap between the goals of BiasEnsemble and our method resulted in smaller improvement of our method on BiasEnsemble-based baselines.
Table 5: Results on Colored MNIST dataset show that combining DiD not only boosts the performance of existing debiasing methods but also achieves the best performances. The accuracy of BN samples is marked as ’-’ in LMLP and HMHP distribution for there is no BN sample within the dataset according to our evaluation setting in Appendix D.
@@
Table 6: Results on Corrupted CIFAR10 dataset show that combining DiD not only boosts the performance of existing debiasing methods but also achieves the best performances. The accuracy of BN samples is marked as ’-’ in LMLP and HMHP distribution for there is no BN sample within the dataset according to our evaluation setting in Appendix D.
Distr. Algorithm Accuracy BA acc BC acc BN acc Avg acc LMLP ERM 80.40 ± 0.81 62.50 ± 0.15 - 64.29 ± 0.06 LfF 59.13 ± 0.68 55.03 ± 0.04 - 55.44 ± 0.09 + DiD 69.47 ± 0.96 62.04 ± 0.21 - 62.78 ± 0.10 BEL 70.87 ± 1.30 52.10 ± 0.30 - 53.98 ± 0.40 + DiD 63.23 ± 2.10 53.21 ± 0.20 - 54.21 ± 0.38 DisEnt 61.58 ± 0.57 55.45 ± 0.23 - 56.06 ± 0.17 + DiD 72.23 ± 0.74 60.84 ± 0.40 - 61.98 ± 0.30 BED 62.73 ± 0.61 56.59 ± 0.08 - 57.20 ± 0.13 + DiD 65.98 ± 0.40 60.92 ± 0.20 - 61.42 ± 0.21 HMLP ERM 84.67 ± 0.64 55.85 ± 0.17 65.75 ± 0.00 65.05 ± 0.13 LfF 73.33 ± 1.67 47.70 ± 0.58 54.58 ± 0.49 54.15 ± 0.41 + DiD 78.67 ± 2.14 54.81 ± 2.26 63.71 ± 2.69 63.06 ± 2.63 BEL 70.33 ± 2.19 50.96 ± 2.35 54.14 ± 0.25 54.02 ± 0.36 + DiD 68.80 ± 0.88 50.20 ± 0.79 54.39 ± 0.18 54.15 ± 0.15 DisEnt 61.67 ± 1.67 52.48 ± 0.56 54.65 ± 0.56 54.53 ± 0.49 + DiD 73.67 ± 2.64 55.26 ± 0.93 62.11 ± 0.17 61.61 ± 0.13 BED 75.33 ± 5.21 49.15 ± 1.54 56.86 ± 0.30 56.35 ± 0.35 + DiD 78.40 ± 1.00 54.09 ± 1.07 62.05 ± 0.34 61.50 ± 0.38 HMHP ERM 89.97 ± 0.34 29.37 ± 0.30 - 35.43 ± 0.24 LfF 72.70 ± 0.81 35.30 ± 0.33 - 39.04 ± 0.33 + DiD 82.07 ± 1.09 37.05 ± 0.31 - 41.55 ± 0.19 BEL 82.73 ± 0.92 31.48 ± 0.82 - 36.61 ± 0.65 + DiD 78.30 ± 0.47 32.90 ± 1.79 - 37.44 ± 1.61 DisEnt 70.77 ± 2.27 36.04 ± 0.62 - 39.51 ± 0.36 + DiD 76.60 ± 0.70 39.05 ± 0.35 - 42.80 ± 0.25 BED 78.60 ± 1.56 34.20 ± 0.43 - 38.64 ± 0.38 + DiD 78.70 ± 1.47 37.72 ± 0.96 - 41.82 ± 0.91
E.2 Results on more existing debiasing methods
In Table 8, we show the results on more existing debiasing methods to show the generality of DiD. The results show that our approach is consistently effective on DPR and DeNetDM.
Table 8: DiD still effectively boosts the performance of even very recent methods. This further demonstrated the adaptability of our approach. The experiments are conducted based on Corrupted CIFAR10.
LMLP HMLP HMHP
Algorithm BC acc Avg acc BC acc Avg acc BC acc Avg acc
DPR 51.54 54.25 43.67 44.67 25.92 31.77
+ DiD +6.97 +4.36 +5.44 +13.16 +2.61 +2.47
DeNetDM 60.18 61.98 49.67 62.11 24.48 31.25
+ DiD +1.93 +2.00 +3.66 +1.30 +3.21 +2.99
E.3 Debias on unbiased datasets
As we do not know how biased or is the training data biased at all in real-world scenarios, it is important to evaluate the performance of debiasing methods on unbiased training data to ensure that they do not cause severe performance degradation, if not improving the performance. As shown in Table 11, existing methods perform poorly on unbiased training data, causing severe performance degradation, yet our approach greatly boosts their performances.
E.4 Experiments on multi-bias settings
The existence of multiple biases is another challenge in debiasing in the real world [22]. We further propose multiple biases with different magnitude and prevalence (HMLP + LMLP) based on Corrupted CIFAR10 benchmark, containing 10 target features and 20 spurious features. As shown in Table 9, our approach is consistently effective on multiple bias settings, debiasing multiple biases at the same time.
Interestingly, no signs of the whac-a-mole phenomenon is observed. We suspect that the phenomenon might occur only between two extremely strong biases, as assumed in the original paper.
Table 9: Experiments on the multiple bias setting with LMLP and HMLP combined has demonstrated consistent effectiveness of our approach in handling multiple biases.
Algorithm LMLP BC HMLP BC Avg
LfF 47.09 48.56 48.27
+ DiD +9.40 +10.66 +9.74
DisEnt 50.82 49.15 51.09
+ DiD +6.70 +7.01 +7.00
BEL 53.66 56.33 54.72
+ DiD +2.23 +2.89 +2.69
BED 56.26 54.66 56.54
+ DiD +1.02 +1.20 +0.94
E.5 Application of DiD on the bias detection task